OpenStudy (maheshmeghwal9):

Two bodies fall freely from different heights & reach the ground simultaneously. The time of descent for the first body is t1=2sec. & for the second, t2=1sec. At what height was the first body situated when the other began to fall????? Answer: --------- 14.7 meters.

OpenStudy (anonymous):

first find out the original ht from which both b1 and b2 fall using s = ut + 0.5at^2. (taking g = 10m/s^2) u should get s1 = 20m and s2 = 5m.

OpenStudy (anonymous):

now, before b2 starts falling, say b1 covers a distance x. so it need s to cover the remaining distance, ie, 20-x in 1 sec (time taken for b2 to reach the ground). so that both of them reach the ground simultaneously. the total time taken by b1 to cover the 20m is 2sec. x is the distance it covers in 2-1 = 1sec. use x = ut+0.5at^2 to get x. then 20-x will give u the answer.

OpenStudy (maheshmeghwal9):

oh ok thanx:)

OpenStudy (anonymous):

np :)